OpenStudy (anonymous):

For a satellite to be in a circular orbit 760kmkm above the surface of the earth. What is the period of the orbit (in hours)?

OpenStudy (anonymous):

The equation you'll to consider is the following:\[\frac{ T ^{2} }{ r ^{3} }=\frac{ 4\pi ^{2} }{ GM _{E} }\]where T is the orbital period of the orbiting satellite; r is the mean orbital radius; G is the gravitational constant; and ME is the mass of the Earth. Note two things: 1. The period will be in seconds, so you'll need to convert. 2. The mean orbital radius is measured from the center of the Earth.

OpenStudy (mrnood):

Just for your info @tornadosiren - the formula given above can be derived from some basic principles: Gravitational force between two bodies = \[F= \frac{ GMm }{ r ^{2} }\] And for a circular path the centripetal force required is \[mr \omega ^{2}\] If you equate those two forces: \[\omega ^{2}=\frac{ GM }{ r ^{3} }\] Since \[T=\frac{ 2\pi }{ \omega } \] This gives the equation above